Find an equation of the tangent plane to the surface at the given point. f(x, y) = x2 − 2xy + y2, (1, 5, 16) with maple lab please


1
Expert's answer
2022-01-06T16:54:01-0500

Solution:


The equation of the tangent plane to the surface at the point "(x_o,y_o,z_o)" on the surface, is:



"z-z_o=f_x(x_o,y_o)(x-x_o)+f_y(x_o,y_o)(y-y_o)"



In the equation, the coefficients are the partial derivatives of the function.


By using the two-dimensional surface "z=f(x,y)", we can evaluate the partial derivatives of the function with respect to x and y at the given point.


Given surface is "z=f(x,y)=x^2-2xy+y^2" at the given point (1, 5, 16).


First, we find the partial derivatives of the surface at the given point:


"\frac{∂f}{∂x}=\frac{∂}{∂x}(x^2-2xy+y^2)"


"\frac{∂}{∂x}(x^2)-\frac{∂}{∂x}(2x)+\frac{∂}{∂x}(y^2)"


"f_x(x,y)=2x-2y"

"f_x(1,5)=2(1)-2(5)=2-10"

"f_x(1,5)=-8"


"\frac{∂f}{∂y}=\frac{∂}{∂y}(x^2-2xy+y^2)=\frac{∂}{∂y}(x^2)-\frac{∂}{∂y}(2xy)+\frac{∂}{∂y}(y^2)"


"f_y(x,y)=2y-2x"

"f_y(1,5)=2(5)-2(1)=10-2"

"f_y(1,5)=8"


Let us find an equation of the tangent plane to the surface at the given point:



"z-z_o=f_x(x_o,y_o)(x-x_o)+f_y(x_o,y_o)(y-y_o)"


"z-16=(-8)(x-1)+(8)(y-5)"


"z-16=-8x+8y-32"


"z=-8x+8y-32+16"


"z=-8x+8y-16"


An equation of the tangent plane to the surface at the given point is 

"z=-8x+8y-16"


Maple code:


TangentPlane(x^2 − 2xy + y^2,x=1,y=5,z=16)


output:


"8x-8y+z=-16"